On January 1, 2021, a company issues $740,000 of 8% bonds, due in nine years, with interest payable semiannually on June 30 and December 31 each year. Assuming the market interest rate on the issue date is 9%, the bonds will issue at $695,008.

Required:
1. Fill in the blanks in the amortization schedule below: (Round your answers to the nearest dollar amount. Enter all amounts as positive values.)

Date Cash Paid Interest Expense Change in Carrying Value Carrying Value
01/01/2021
06/30/2021
12/31/2021

2. Record the bond issue on January 1, 2021, and the first two semi-annual interest payments on June 30, 2021, and December 31, 2021. (If no entry is required for a particular transaction/event, select "No Journal Entry Required" in the first account field. Round your answers to the nearest dollar amount.)
Answer #1
1
Date Cash Paid Interest Expense Change in Carrying Value Carrying Value
01/01/2021 695008
06/30/2021 29600 31275 1675 696683
12/31/2021 29600 31351 1751 698434
2
Debit Credit
January 1, 2021 Cash 695008
Discount on Bonds payable 44992
      Bonds payable 740000
June 30, 2021 Interest expense 31275 =695008*9%/2
     Discount on Bonds payable 1675
     Cash 29600 =740000*8%/2
December 31, 2021 Interest expense 31351 =696683*9%/2
     Discount on Bonds payable 1751
     Cash 29600